OpenStudy (anonymous):

How do you solve... Two cars are 400 mi apart. Both start at the same time and travel toward one another. They meet 4 hr later. If the rate of one car is 20 mph faster than the other, what is the rate of each car?

ganeshie8 (ganeshie8):

say, the speed of slow car = x then, speed of fast car = x+20

ganeshie8 (ganeshie8):

in 4 hours, they both combinedly travelled 400mi

ganeshie8 (ganeshie8):

use below formula :- distance = speed * time

ganeshie8 (ganeshie8):

400 = (x+x+20)*4

OpenStudy (anonymous):

40?

ganeshie8 (ganeshie8):

yup, speed of slow car, x = 40 speed of fast car = x+20 = 40+20 = 60

OpenStudy (anonymous):

Thank you! :)

ganeshie8 (ganeshie8):

np :)

ganeshie8 (ganeshie8):

If it helps, u may think like this :- we have added speeds x+x+20, cuz speeds are helping each other - to quickly meet each other